OpenStudy (anonymous):

Which of the following is the slope-intercept form of the equation 8x + y = -1

OpenStudy (anonymous):

A. y = -8x − 1B. y = -8x + 1C. y = 8x − 1D. y = 8x + 1 the answers

OpenStudy (anonymous):

A

OpenStudy (anonymous):

as you put 8x to the other side so it becomes negative and then you have y= -8x -1

OpenStudy (anonymous):

its a -8x-1

OpenStudy (anonymous):

-8 is the slope, -1 is the intercept

OpenStudy (maheshmeghwal9):

A. as \[8x+y=-1\]then \[y=-1-8x\]

OpenStudy (maheshmeghwal9):

Am I correct?

OpenStudy (anonymous):

yes

OpenStudy (maheshmeghwal9):

So didvu get it?:)

OpenStudy (anonymous):

yes

OpenStudy (maheshmeghwal9):

K!